The table below describes results from group of eight births from eight different sets of parents. The random variable x represents the number of girls among eight children.Use the table to answer questions (a) to (e) below. 3. Number of Girls x Pk) 0.004 0.109 0.219 0.273 0.219 0.109 6 7 0.004 a Find the probability of getting exactly one girs in eight births b. Find the probability of getting one or fewer girls in eight births. c. Which probability is relevant for determining whether one is a significantly low number of girls in eight births: the result from part (a) or part (b)? Explain. d. Is one a significantly low number of girls in eight births? Why or why not?
Answer #1

a) P(One girl in eight births) = 0.031

b) P(One or fewer girls in eight births) = P(0 girls) + P(1 girl) = 0.004 + 0.031 = 0.035

c) The result from part b) is relevant because to check the significance of one birth, we need to look for the probability of one or fewer births.

d) Yes, because the probability of one or fewer girls is less than 0.05